OpenStudy (anonymous):

PLEASE ANSWER: A ribbon costs $2 for each yard. Give the cost of each length. A. 2/3 yard B.1 1/2 feet C. 8 inches D. 1 3/4

OpenStudy (anonymous):

a. 2/3 * $2 = 1.33 b. 1.5 ft * 1/12 * $2 = .25 c. 8 in * 1/36 * 2 = .44 d. what is sthe units